What is the LCM of 6, 7, and 9?
Answer by rothauserc(4718)

Write prime factorization for each number
:
6 = 3 * 2
7 = 7
9 = 3 * 3
:
make a list of the most occurrences of each prime
:
2 --》1
3 --》2
7 --》1
:
Therefore LCM = 2 * 3 * 3 * 7 = 126
: